OpenStudy (anonymous):

If there is 360 grams of radioactive material with a half-life of 8 hours, how much of the radioactive material will be left after 32 hours and is the radioactive decay modeled by a linear function or an exponential function? A) 22.5 grams; linear B) 22.5 grams; exponential C) 45 grams; linear D) 45 grams; exponential

OpenStudy (anonymous):

@Aaron96

OpenStudy (shamim):

Required formula is N=N0*e^(-lamda*t)

OpenStudy (anonymous):

?

OpenStudy (aaron96):

360/4 = 90 considering it is half life then it would be 45

OpenStudy (shamim):

Here N0=360gram=initial amount of radioactive materials

OpenStudy (anonymous):

ok

OpenStudy (shamim):

N=remaining amount of radioactive matetials after 32 hrs

OpenStudy (shamim):

U hv to find out N=?

OpenStudy (shamim):

e means exponential

OpenStudy (anonymous):

im lost

OpenStudy (shamim):

Ok

OpenStudy (shamim):

U know an equation Lamda=0.693/half life

OpenStudy (anonymous):

yea

OpenStudy (shamim):

Lamda=?

OpenStudy (shamim):

Half life is given

OpenStudy (shamim):

Half life=8 hrs

OpenStudy (jclark):

360/2=180/2=90/2=45/2=22.5

OpenStudy (bearjeau):

Just divide by 2 every 8 hours